how i get 50 by the use 5 digit in the use of 1,3,5,7,9,11,13,15,17,19
Answer by Edwin McCravy(20065)

There are many ways:
(3×13)+17-(1+5) = 50
1×(3+11+17+19) = 50
3×(1+5+7)+11 = 50
5×(1+3)+13+17 = 50
7×(3+1)+9+13 = 50
1×(19+15+13)+3 = 50
1×(3+11)+17+19 = 50
3×(9+5)+7+1 = 50
Edwin